a motorist drove 150 miles on country roads before driving 50 miles on mountain roads. the rate of speed on the country roads was 3 times the rate on mountain roads. the time spent traveling the 200 miles was 5 hours. find the rate of the motorist on country roads?
i let x=rate of the motorist on mountain road
and 3x rate of the motirist on country roads
i set up the equation x/200 - 3x/200 = 5

Answer by stanbon(75887)

a motorist drove 150 miles on country roads before driving 50 miles on mountain roads.
the rate of speed on the country roads was 3 times the rate on mountain roads.
the time spent traveling the 200 miles was 5 hours.
find the rate of the motorist on country roads?
---
Country Road DATA:
distance = 150 miles ; rate = 3x mph ; time = d/r = 150/(3x) = 50/x hrs.
--------------------------------
Mountain Road DATA:
distance = 50 miles ; rate = x mph ; time = d/r = 50/x hrs
---------------------------------
Equation:
time + time = 5 hrs
2(50/x) = 5
100/x = 5
5x = 100
x = 20 mph on mountain Roads
3x = 60 mph on country Roads
=====================================
Cheers,
Stan H.
=====================================

i let x=rate of the motorist on mountain road
and 3x rate of the motirist on country roads
i set up the equation x/200 - 3x/200 = 5
Comment: x/200 is time/distance = rate
But 5 is time
You need an equation that says time + time = time.